OpenStudy (k8lyn911):

Calculus III Homework Help: If a vector has direction angles A = π/4 and B = π/3, find the third direction angle C.

OpenStudy (danjs):

I think the sum of the squared of the 3 direction angles is identically 1

OpenStudy (danjs):

squares

OpenStudy (danjs):

direction cosine, sorry

OpenStudy (danjs):

Barely used that except for that one small section in calc 3 if i remember, not too important, but good to know

OpenStudy (k8lyn911):

Okay, I think I've gotten it. Thank you so much for your help! :)

OpenStudy (danjs):

cos^2(C) = 1/4 cos(C) = 1/2 pi/3

OpenStudy (danjs):

welcome